### v6

@@ 4.1 Preliminary lemmas
$I_{1}=-p\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s.$ (4.36)
Regarding $I_{2}$, we can deduce that
$I_{2}$ $\leqslant\,c\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\big{(}\|{\xi_{h}(s)}\|_{L^{2}}^{2}+\|{\eta_{h}(s)}\|_{L^{2}}^{2}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{2}\big{)}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\,\mathbb{E}\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{2}}^{p}+\|{\eta_{h}(s)}\|_{L^{2}}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
=$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{=}}c\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.37)
where (i) follows from Young’s inequality, (ii) uses the determinism of $G$ and $G_{h}$, and (iii) is justified by the inequalities (4.25) and (4.27). For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.2, a straightforward calculation yields
$I_{3}$ $\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s$
$\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s$
$\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s.$
@@
Using (4.30), we further refine this bound to:
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s.$
Recalling the definitions in Eqs. 4.32 and 4.33 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
@@
which, in conjunction with Eq. 4.25, Eq. 4.26, and Eq. 4.27, yields
$\mathbb{E}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
@@
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.36, 4.37, 4.38, 4.39 and 4.41, we obtain the following inequality:
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s$
@@
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.11, and let $z$ be defined by Eq. 4.32. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for $y$ and $z$:
@@
Let $G_{h}$ be defined by Eq. 2.12 and let $z_{h}$ be defined by Eq. 4.33. We adopt the same notation introduced in the proof of Theorem 3.2:
$e_{h}$ $:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$
$\eta_{h}$ $:=e_{h}-\xi_{h}.$
It is straightforward to confirm the following decomposition for $y-y_{h}$:
$y-y_{h}=(z-P_{h}z)+(G-G_{h})-(\xi_{h}+\eta_{h}).$ (4.45)
@@
Estimate (4.46) is a direct consequence of (4.3) combined with the regularity property (4.43), and estimate (4.47) follows from (4.6) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Analogous to the estimate Eq. 4.27, we employ Lemma 4.5, along with (4.4) and the regularity assertion (4.43), to bound $\xi_{h}$ as follows:
@@
Adopting the reasoning from Theorem 3.2, we establish the following bound for $\eta_{h}$:
@@
In light of the decomposition Eq. 4.45, combining Eqs. 4.48, 4.49, 4.47 and 4.46 leads to the error estimate
@@
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded w.r.t~{}$h$ for all $p,q\in[2,\infty)$}.$ (4.51)
@@
Using the equality Eq. 4.35, along with the definitions of the operators $S_{0,h}$ and $S_{1,h}$, we deduce the following decomposition:
$\eta_{h}=I_{1}+I_{2}+I_{3},$ (4.53)
@@
$I_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$I_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$I_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $I_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
@@
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.5; and the final inequality is justified by Eqs. 4.49, 4.47 and 4.48. Turning to $I_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
@@
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.5 and Eq. 4.2; the third inequality uses Hölder’s ineqality; and the last inequality uses Eqs. 4.42, 4.50 and 4.51. For $I_{3}$ we demonstrate the following bound:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
@@
[…]
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.6, along with Eqs. 4.2 and 4.10. Given that the decomposition Eq. 4.53 implies
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
$\qquad{}+\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})},$
combining the above bounds for $I_{1}$, $I_{2}$ and $I_{3}$ yields Eq. 4.52.
Step 3. Let us proceed to prove the error estimate Eq. 3.7. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. We start by establishing the following error estimate:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{6})}+{}$
[…]
Employing this error estimate and following the proof methodology for bounding $I_{3}$ in Step 2, we further obtain an enhanced bound for $I_{3}$:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$
Integrating this refined bound with those previously derived bounds for $I_{1}$ and $I_{2}$ in Step 2, and considering the decomposition in Eq. 4.53, we conclude that
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$ (4.54)
Furthermore, we infer that
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+{}$
[…]
Consequently, combining these estimates directly yields the desired error estimate Eq. 3.7.
@@
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $I_{1}$ and $I_{2}$ from (4.53), in conjunction with Lemma 4.10 and the continuous embedding of $\dot{H}_{h}^{\varepsilon,q}$ into $L^{q}$ for any $\varepsilon>0$, we deduce
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
@@
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
@@
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
@@
Consequently, we obtain the following bound for $I_{1}+I_{2}$:
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $I_{3}$ from Eq. 4.53 as follows:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
@@
where the first inequality is justified by Lemma 4.6 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.2; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.10. Combining the above bounds for $I_{1}+I_{2}$ and $I_{3}$, we arrive at
$\lVert{I_{1}+I_{2}+I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
@@
Furthermore, combining Eqs. 4.47 and 4.55 leads to the following bound:
@@
Given that the decomposition Eq. 4.45 implies
@@
the desired error estimate Eq. 3.8 is derived by combining the above bounds. This concludes the proof of Theorem 3.3.

### v7

@@ 4.1 Preliminary Lemmas
$I_{1}=-p\mathbb{E}\biggl{[}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s\biggr{]}.$ (4.39)
Regarding $I_{2}$, by Hölder’s inequality, Young’s inequality, and the inequalities Eq. 4.27 and Eq. 4.29, we can deduce that
$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.40)
For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.18, a straightforward calculation yields
$I_{3}$ $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s\bigg{]}$
$\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s\bigg{]}.$
@@
Using (4.32) and (4.33), we further simplifies this bound to:
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
@@
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}.$
Recalling the definitions in Eqs. 4.34 and 4.35 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
@@
which, in conjunction with Eqs. 4.27, 4.28 and 4.29, yields
$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
@@
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.39, 4.40, 4.41, 4.42 and 4.44, we obtain the following inequality:
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\bigg{[}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s\bigg{]}$
@@
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.10, and let $z$ be defined by Eq. 4.34. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for the processes $y$ and $z$:
@@
Let $G_{h}$ be defined by Eq. 4.36. We adopt the same notation as introduced in the proof of Theorem 3.2:
$z_{h}$ $:=y_{h}-G_{h},$ $e_{h}:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$ $\eta_{h}:=e_{h}-\xi_{h}.$
@@
Estimate (4.48) is a direct consequence of (4.19) combined with the regularity property (4.46). Estimate (4.49) follows from (4.22) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. In analogy to the estimate (4.29), we utilize Lemma 4.7, along with (4.20) and the regularity assertion (4.46), to derive the following bound for $\xi_{h}$:
@@
Following the reasoning from Theorem 3.2, we obtain the following estimate for $\eta_{h}$:
@@
In light of the decomposition Eq. 4.43, combining Eqs. 4.50, 4.51, 4.49 and 4.48 leads to the error estimate
@@
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded with respect to $h$ for all $p,q\in[2,\infty)$}.$ (4.53)
@@
Using the equality Eq. 4.38, along with the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.24 and 4.25, we deduce the following decomposition:
$\eta_{h}=II_{1}+II_{2}+II_{3},$ (4.55)
@@
$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$II_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$II_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $II_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{II_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
@@
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.7; and the final inequality is justified by Eqs. 4.51, 4.49 and 4.50. Turning to $II_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
@@
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.7 and Eq. 4.18; the third inequality uses Hölder’s inequality; and the last inequality uses Eqs. 4.45, 4.52 and 4.53. For $II_{3}$ we establish the following bound:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\lVert{S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
@@
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.8, along with Eqs. 4.18 and 4.4. In view of the decomposition Eq. 4.55, combining the above bounds for $II_{1}$, $II_{2}$, and $II_{3}$ yields Eq. 4.54.
Step 3. We proceed to establish the error estimate (3.7). Considering the decomposition of $y_{h}-y$ given in (4.43), we integrate the inequalities (4.48), (4.49), (4.50), and (4.54) to validate (3.7) for the specific case where $p\in(2,\infty)$ and $q=6$. Building upon this outcome and employing a similar bounding technique as utilized for $II_{3}$ in Step 2, we obtain an enhanced bound for $II_{3}$:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$
Integrating this refined bound with the previously established bounds for $II_{1}$ and $II_{2}$ from Step 2, and incorporating the decomposition (4.55), we arrive at the following conclusion:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.56)
Ultimately, revisiting the decomposition (4.43) and synthesizing the inequalities (4.48), (4.49), (4.50), and (4.56), we affirm the validity of the error estimate (3.7) for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
@@
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $II_{1}$ and $II_{2}$ from (4.55), in conjunction with Lemma 4.9 and the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
@@
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
@@
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
@@
Consequently, we obtain the following bound for $II_{1}+II_{2}$:
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $II_{3}$ from Eq. 4.55 as follows:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
@@
where the first inequality is justified by Lemma 4.8 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.18; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.4. Combining the above bounds for $II_{1}+II_{2}$ and $II_{3}$, we arrive at
$\lVert{II_{1}+II_{2}+II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
@@
Furthermore, by combining Eqs. 4.49 and 4.57, we obtain the following bound:
@@
Since the decomposition Eq. 4.43 implies
@@
the desired error estimate Eq. 3.8 follows by combining the above bounds. This concludes the proof of Theorem 3.3.